OpenStudy (anonymous):

How can you tell when a quadratic equation has no real solutions? A. when the radicand is negative B. when b in the quadratic formula is greater than the radicand C. when the radicand equals zero D. when the radicand is not a perfect square I believe the answer is A? is that right?

OpenStudy (mertsj):

A

OpenStudy (campbell_st):

its A using the discriminant \[b^2 - 4ac\] < 0 no real solutions the radical will be negative

OpenStudy (experimentx):

whats radicand anyway??

OpenStudy (experimentx):

i go with campbell

OpenStudy (campbell_st):

the square root symbol

OpenStudy (mertsj):

b^2-4ac

OpenStudy (mertsj):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Thanks!! :) I also have this one... For which value of x does the graph of y = 2x2 − 7x + 6 cross the x-axis? A. −3/2 B. −2/3 C. 2 D. 3 I think its C.

OpenStudy (mertsj):

\[b^2-4ac\] is the discriminant

OpenStudy (mertsj):

You are correct.

OpenStudy (campbell_st):

you can substitute each value to find y = 0 or factorise (2x - 3)(x - 2) = 0 x = 3/2 and 2

OpenStudy (anonymous):

Thanks so much, here's another one i'm stuck on... What are the approximate solutions of 4x2 + 3 = −12x to the nearest hundredth? A. x ≈ −3.23 and x ≈ 0.23 B. x ≈ −2.72 and x ≈ −0.28 C. x ≈ 0.28 and x ≈ 2.72 D. x ≈ −0.23 and x ≈ 3.23 I think its C, but I'm not sure??

OpenStudy (campbell_st):

put it in standard form 4x^2 + 12x + 3 = 0 use the general quadratic formula \[x = (-b \pm \sqrt{b^2 - 4ac})/2a\] in your question a = 4, b = 12 and c = 3

OpenStudy (mertsj):

C is correct.

OpenStudy (anonymous):

thanks so much!! :)

OpenStudy (mertsj):

No. It's B

OpenStudy (mertsj):

Both roots are negative.

OpenStudy (anonymous):

B? Really? are you sure?

OpenStudy (anonymous):

Okay, i'll trust you! :)

OpenStudy (mertsj):

\[x=\frac{-12\pm \sqrt{96}}{8}\]

OpenStudy (anonymous):

I have one more question that's really confusing me, would you mind sticking around for one more?

OpenStudy (mertsj):

\[x=\frac{-12+9.80}{8} or x=\frac{-12-9.80}{8}\]

OpenStudy (mertsj):

ok

OpenStudy (anonymous):

*note* don't ask me why the numbers are spreed out like that, I just copied and pasted it onto here, that how it looks on my paper. Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (mertsj):

b^2-4ac

OpenStudy (anonymous):

is that the answer?

OpenStudy (mertsj):

Well which part would you choose? -b? 2a????

OpenStudy (mertsj):

\[b^2-4ac=7^2-4(2)(3)=49-24=25\]

OpenStudy (anonymous):

I don't know... This is really confusing me for some reason...

OpenStudy (anonymous):

so the answer is b^2-4ac?

OpenStudy (anonymous):

I'm still confused... :/

OpenStudy (anonymous):

or is the answer this: b² - 4ac = (7)² - 4(2)(3) = 49 - 24 = 25.

OpenStudy (mertsj):

The discriminant, which is b^2-4ac, is the part of the quadratic formula which tells you about the roots.

OpenStudy (mertsj):

Do you understand that the question has two parts? READ THE QUESTION!!!

OpenStudy (mertsj):

PART 1: Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring?

OpenStudy (mertsj):

The answer to that is :b^2-4ac

OpenStudy (mertsj):

PART 2: Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (mertsj):

The answer to that is 25

OpenStudy (anonymous):

ohhh okay, I understand now

OpenStudy (anonymous):

Thank you!!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

I just wasn't looking at the problem correctly, sorry for the confusion and thanks again for breaking it down for me